# New AGN diagnostic diagrams based on the [OIII] auroral line

### v1

@@
The James Webb Space Telescope (JWST) is revolutionizing our understanding of black hole formation and growth in the early Universe. However, JWST has also revealed that some of the classical diagnostics, such as the BPT diagrams and X-ray emission, often fail to identify narrow line TypeII active galactic nuclei (AGN) at high redshift. Here we present three new rest-frame optical diagnostic diagrams leveraging the [OIII]auroral line, which has been detected in several JWST spectra. Specifically, we show that high values of the [OIII]Hratio provide a sufficient (but not necessary) condition to identify the presence of an AGN, both based on empirical calibrations (using local and high-redshift sources) and a broad range of photoionization models. These diagnostics are able to separate much of the AGN population from Star Forming Galaxies (SFGs). This is because the average energy of AGNās ionizing photons is higher than that of young stars in SFGs, hence AGN can more efficiently heat the gas, therefore boosting the [OIII]line. We also found independent indications of AGN activity in some high-redshift sources that were not previously identified as AGN with the traditional diagnostics diagrams, but that are placed in the AGN region of the diagnostic presented in this work. We note that, conversely, low values of [OIII]Hcan be associated either with SFGs or AGN excitation. We note that the fact that strong auroral lines are often associated with AGN does not imply that they cannot be used for direct metallicity measurements (provided that proper ionization corrections are applied), but it does affect the calibration of strong line metallicity diagnostics.
@@
Thanks to the successful launch of the James Webb Space Telescope (JWST; Gardner etĀ al., 2023; Rigby etĀ al., 2023), we can now investigate with high resolution and unprecedented sensitivity both the photometric and spectroscopic properties of galaxies up to (Curtis-Lake etĀ al., 2022; Robertson etĀ al., 2022). Within this context, recent studies, exploiting different kinds of JWST data, have revealed a large population of AGN at high redshift (Kocevski etĀ al., 2023; Ćbler etĀ al., 2023a; Matthee etĀ al., 2023; Maiolino etĀ al., 2024, 2023a; Greene etĀ al., 2023; BogdĆ”n etĀ al., 2023; Goulding etĀ al., 2023; Kokorev etĀ al., 2023; Furtak etĀ al., 2023; Juodžbalis etĀ al., 2024; Scholtz etĀ al., 2023a; Chisholm etĀ al., 2024), suggesting a higher AGN fraction at than previously expected. For instance, Yang etĀ al. (2023), taking advantage of the JWST-MIRI photometry of the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein etĀ al., 2022), investigated the AGN population using spectral energy distribution (SED) modelling, and found a black-hole accretion rate density (BHARD) at dex higher than what was expected from previous X-ray studies (Vito etĀ al., 2016, 2018). Also Maiolino etĀ al. (2023a) and Harikane etĀ al. (2023), by selecting broad-line AGN (BLAGN, or TypeI) among NIRSpec spectra, have found significant AGN excess at with respect to the AGN luminosity functions derived using X-ray data (Giallongo etĀ al., 2019).
@@
The X-ray emission, produced in AGN by inverse Compton scattering of the UV photons coming from the SMBH accretion disc, has been traditionally used to select AGN. Most of the high-z AGN selected using JWST spectroscopy lie in fields covered by some of the deepest X-ray observations ever performed, like the 7Ms Chandra Deep Field South (CDFS) (Luo etĀ al., 2017), Chandra Deep Field North (CDFN) (Xue etĀ al., 2016) and AEGIS-XD (Nandra etĀ al., 2015). However, most of these newly discovered high-z AGN lack any X-ray emission (e.g. Maiolino et al. 2024 in prep. Mazzolari et al. 2024 in prep. Scholtz etĀ al., 2023a; Ćbler etĀ al., 2023a), with only very few exceptions (Goulding etĀ al., 2023; Kovacs etĀ al., 2024). This makes the picture even more difficult to interpret and suggests different scenarios such as Compton thick obscuration (Gilli etĀ al., 2022) or intrinsic X-ray weakness (Simmonds etĀ al., 2016).
@@
The difficulties in selecting high-z AGN are even more important for TypeII or narrow-line AGN (NLAGN). Different observational works have already demonstrated that some of the traditional well-known AGN emission-line diagnostic diagrams, like the BPT (Baldwin etĀ al., 1981) diagram, are no longer useful for the high-z Universe (Maiolino etĀ al., 2023a; Scholtz etĀ al., 2023a; Harikane etĀ al., 2023; Kocevski etĀ al., 2023; Ćbler etĀ al., 2023a; Groves etĀ al., 2006). Indeed, at high-z, the environments are systematically more metal poor and the stellar populations are generally younger, increasing the ionization parameter () of star-forming galaxies (SFG). The latter effect, on the one hand, makes SFGs move towards the AGN locus on the BPT; on the other hand the lower metallicity of the Narrow Line Region (NLR) of AGN makes its line ratios move towards and overlap with the star formation locus on these traditional diagnostic diagrams. The detection of high-ionization emission lines (like NāVā, [NeāIV], [NeāV]) can still be considered as a safe tracer for the presence of an AGN, but their detection is still difficult even in the deepest JWST spectroscopic surveys like the JWST Advanced Deep Extragalactic Survey (JADES, Bunker etĀ al., 2023; Scholtz etĀ al., 2023a).
@@
However, Brinchmann (2023) first noticed that the anomalously high [OIII]emission in some of the early NIRSpec spectra at z8 could be indicating the presence of an AGN. More recently, Ćbler etĀ al. (2023b) have suggested that a strong ratio between this line and the Hline could be a possible indicator for the presence of an AGN. In their work, by using JWST/NIRSpec Integral Field Spectroscopy (IFS) observations, they noticed a spatial correspondence between the AGN broad Hemission and the [OIII]line emission. They also mention the possibility that the enhanced [OIII]emission in correspondence of the AGN could be related to higher ISM temperatures driven by the AGN activity.
@@ 2.1 Low Redshift Samples
The sample of low-z SFGs is constituted by both normal SFGs and also by the so-called local analogues of high-z galaxies, i.e. local galaxies that show intrinsic properties similar to those observed in high-z galaxies, in terms of metallicities and emission lines ratios (e.g. Izotov etĀ al., 2018, 2019, 2021). The Sloan Digital Sky Survey (SDSS) DR7 (Abazajian etĀ al., 2009) gives the most numerous sample, mostly made of galaxies. Starting from the whole catalog provided by MPA/JHU, we considered only sources with a in all the lines involved in the diagnostic diagrams we are going to present: [OIII], [OIII], [NeāIII], [OII], and H. Then we distinguished between AGN and SFGs using the BPT diagram (Baldwin etĀ al., 1981), taking the AGN demarcation line provided by Kewley etĀ al. (2001) and after considering a further cut in SNR also in the H, Hand [NāII]Ā lines. Furthermore, we exclude from the selection AGN classified in the SDSS catalog as BLAGN, since the catalog does not provide separately the NL and BL components of the Balmer lines and our diagnostics rely on the NL emission only. We also correct all the lines for dust attenuation using the Calzetti etĀ al. (2000) attenuation law. The final SDSS SFG (AGN) sample considered in this work contains () sources.
@@
The last sample of local AGN is taken from the MOSFIRE Deep Evolution Field (MOSDEF) survey (Kriek etĀ al., 2015), which targeted the rest-frame optical spectra of H-band-selected Galaxies at . We took the full MOSDEF catalogue and selected only those sources with a [OIII]line with a , and then we distinguished between AGN and SFG using the BPT diagram. With these SNR cuts, we could identify only 3 sources, all classified as AGN. This indicates the difficulty in detecting such a faint line further than in the local Universe in the pre-JWST era.
@@ 2.2 High Redshift Samples
The High-z samples of AGN and SFGs only comprise sources observed spectroscopically using JWST, the only instrument with enough resolution, sensitivity and IR-coverage to detect and disentangle the [OIII]line at . We considered the full publicly released sample of sources with a medium resolution (MR, ) spectrum in the CEERS (PID: 1345; Finkelstein etĀ al., 2022) and JADES (PID: 1210; Bunker etĀ al., 2023) surveys. For JADES, we considered the emission line fluxes and the AGN selection described in Scholtz etĀ al. (2023a). Starting from a sample of 110 sources with medium resolution (MR) spectroscopy, a reliable redshift and sufficient wavelength coverage, the authors performed a NLAGN selection using multiple AGN rest-frame UV and optical diagnostic diagrams, finally selecting 28 reliable AGN candidates in the redshift range . Of all the initial JADES sources, only 19 show a [OIII]detection, 7 of which AGN. We followed the same approach for the CEERS program: starting from 313 MR spectra published by the CEERS collaboration, we selected 217 sources with a reliable spectroscopic redshift and we finally selected 45 AGN candidates (Mazzolari et al. 2024 in prep). For this work, we considered 32 sources with the [OIII]line detected among the CEERS sample, 4 of which already classified as AGN. From the same programs, we also collect the BLAGN identified by Maiolino etĀ al. (2023a) and Harikane etĀ al. (2023) at , carefully checking, and in case subtracting, the broad component of the Hline.
@@
The sample of high-z sources presented in Nakajima etĀ al. (2023) includes galaxies from three different early JWST programs (CEERS, ERO, GLASS) at . We selected only the 10 sources with the [OIII]line detected, and we marked as BLAGN those that were selected based on their broad Hline in Harikane etĀ al. (2023). For the latter, we subtract the broad Hcomponent, if present.
We also include in the high-z AGN sample the BLAGN reported in Ćbler etĀ al. (2023a) at , whose JWST/NIRSpec Integral Field Spectrograph (IFS) observation shows very high ionization lines and low metallicity (1/4 solar). We also include both the components of the dual AGN candidate at reported in Ćbler etĀ al. (2023b). For this source, thanks to IFS spectroscopy, the authors were able to study the displacement between the position of the Hbroad line region (BLR) and the strong [OIII]emission line centroid, which were interpreted as the emissions coming from two distinct sources 620pc apart from each other. Interestingly, the authors also found an almost perfect alignment between the [OIII]emission line peak and the peak of the HBLR, suggesting a correlation between the intensity of the [OIII]line and the presence of an AGN.
We also consider the BLAGN reported in Kokorev et al. (2023) at from JWST UNCOVER Treasury survey (Labbe et al., 2021), showing a robust Hbroad component and an unprecedented black hole to host galaxy mass of at least . An even larger black hole to host galaxy mass ratio was more recently found in the BLAGN analyzed in Juodžbalis et al. (2024), representing an extreme example of dormant SMBH at high-z. We also considered the BLAGN, triply imaged and lensed by the galaxy cluster Abell2744-QSO1, reported by Furtak et al. (2023). This source is probably undergoing a phase of rapid SMBH growth and is also heavily obscured, since the authors measured an from the Balmer decrement.
@@
We finally included in the sample also GN-z11, an exceptionally luminous galaxy at z = 10.6. Its JWST spectrum reveals the presence of an AGN through the detection of semi-forbidden lines tracing very high densities (inconsistent with the ISM but typical of the Broad Line Region), other transitions typical of AGN, fast outflows, ionization cones, and a larger Ly halo, consistent with those seen in lower redshift quasars (Maiolino etĀ al., 2024, 2023b; Scholtz etĀ al., 2023b).
@@
Ćbler etĀ al. (2023a) | AGN | 1 | ||
Ćbler etĀ al. (2023b) | AGN | 2 | ||
Juodžbalis et al. (2024) | AGN | 1 | ||
@@ 3 Results
On the right panel, we show the distribution of the photoionization models presented in Sec. 2.3 according to the same line ratios. We see here that SFG models are predicted to occupy a limited region of this diagram, whose upper boundary almost perfectly corresponds with the distribution of SDSS SFG and local analogs, while AGN are expected to cover both the region occupied by SFGs and also the part of the diagnostic covered by the local AGN samples. Therefore, SFG models are not able to reach the part of the diagram occupied by local AGN, as expected also from the distribution of the observational samples. The opposite is not true, since AGN can occupy the region populated by SFGs, as observed for some local and high-z AGN. These considerations remain true even considering the full grid of SFG models computed in Feltre etĀ al. (2016), as shown in Fig. Ā 7 of the appendix.
@@
It is worth noting the position of some high-z sources in this diagnostic. The AGN reported in Kokorev etĀ al. (2023) shows an extremely strong [OIII]line and it is placed in a region that is not covered by any model, which probably suggests a very high electron temperature and , as also pointed out in Kokorev etĀ al. (2023). Another significant high-z AGN falling in the AGN-dominated region of this diagnostic is the source corresponding to the BLR centroid in the dual source described in Ćbler etĀ al. (2023b), where, in the IFS map, the peak of the broad-line emission spatially corresponds to the peak of the [OIII]line emission. The other component, the [OIII]centroid, showing a fainter [OIII]line emission, is close to the border between the local analogues and the local AGN distribution, together with most of the high-z BLAGN. The AGN region is also populated by different AGN selected among the CEERS and JADES MR NIRSPEC spectra, but we also have different sources not selected as AGN but falling in the AGN-dominated region, as pointed out above. We demonstrate in Sec. 4 that, from the stack of their spectra, we have an indication that some of these sources may host a BLAGN too.
@@
Finally, it is worth noting that the effect of dust reddening on the line ratio [OIII]/[OII], whose components can suffer a different amount of attenuation due to their distance in wavelengths, would move data points towards the right. This can potentially shift AGNs into the SFG+AGN region under significant reddening, but not the opposite, i.e. galaxies dominated by star formation would not contaminate the AGN-only region.
@@
The third diagnostic is presented in Fig.Ā 3 and is based on the ratio between the [OIII]/Hcompared to [OIII]/[OIII]. This diagnostic was reported also in Ćbler etĀ al. (2023b). Looking at the distribution of SDSS SFGs and AGN in the left panel, we note that there is more overlap between the two populations with respect to the previous diagnostics, but there is still a region, characterized by high values of [OIII]/Hand high [OIII]/[OIII], that is populated only by local AGN. The local analogs distribution seems again to be an extrapolation of the trend of the SDSS SFG distribution. In this diagram, there are very few high-z galaxies that are in the AGN-only region of the diagnostic, while most of them are shifted towards lower values of [OIII]/[OIII], even in a region that is not covered by the current photoionization models, as it is evident from the right panel. This trend is emblematically represented by the two extreme AGN reported in Kokorev etĀ al. (2023) and in Furtak etĀ al. (2023). Given the strength of their [OIII]line emission, they are placed in the upper-left part of the plot in a region that cannot be reproduced by any model. The same trend is shown by most of the galaxies that were not selected as AGN, but that lie in the AGN region of the two diagnostics presented above. The distribution of the SFG models matches the distribution of the local galaxies, in particular in its boundary with the AGN-only region, while AGN models cover both the region occupied by SFGs and the one populated only by AGN samples.

### v2

@@
The James Webb Space Telescope (JWST) is revolutionizing our understanding of black hole formation and growth in the early Universe. However, JWST has also revealed that some of the classical diagnostics, such as the BPT diagrams and X-ray emission, often fail to identify active galactic nuclei (AGN) at high redshift and low metallicity. Here we present three new rest-frame optical diagnostic diagrams to identify narrow-line Type II AGN, leveraging the [OIII]auroral line, which has been detected in several JWST spectra. Specifically, we show that high values of the [OIII]Hratio provide a sufficient (but not necessary) condition to identify the presence of an AGN, both based on empirical calibrations (using local and high-redshift sources) and a broad range of photoionization models. These diagnostics are able to separate much of the AGN population from Star Forming Galaxies (SFGs). This is because the average energy of AGNās ionizing photons is higher than that of young stars in SFGs, hence AGN can more efficiently heat the gas, therefore boosting the [OIII]line. We also found independent indications of AGN activity in some high-redshift sources (zĀæ4) that were not previously identified as AGN with the traditional diagnostics diagrams, but that are placed in the AGN region of the diagnostic presented in this work. We note that, conversely, low values of [OIII]Hcan be associated either with SFGs or AGN excitation. We note that the fact that strong auroral lines are often associated with AGN does not imply that they cannot be used for direct metallicity measurements (provided that proper ionization corrections are applied), but it does affect the calibration of strong line metallicity diagnostics.
@@
Thanks to the successful launch of the James Webb Space Telescope (JWST; Gardner etĀ al., 2023; Rigby etĀ al., 2023), we can now investigate with high resolution and unprecedented sensitivity both the photometric and spectroscopic properties of galaxies up to (Curtis-Lake etĀ al., 2022; Robertson etĀ al., 2022). Within this context, recent studies, exploiting different kinds of JWST data, have revealed a large population of AGN at high redshift (Kocevski etĀ al., 2023; Ćbler etĀ al., 2023; Matthee etĀ al., 2023; Maiolino etĀ al., 2024b, 2023a; Greene etĀ al., 2023; BogdĆ”n etĀ al., 2023; Goulding etĀ al., 2023; Kokorev etĀ al., 2023; Furtak etĀ al., 2023; Juodžbalis etĀ al., 2024b; Scholtz etĀ al., 2023a; Chisholm etĀ al., 2024), suggesting a higher AGN fraction at than previously expected. For instance, Yang etĀ al. (2023), taking advantage of the JWST-MIRI photometry of the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein etĀ al., 2022), investigated the AGN population using spectral energy distribution (SED) modelling, and found a black-hole accretion rate density (BHARD) at dex higher than what was expected from previous X-ray studies (Vito etĀ al., 2016, 2018). Also Maiolino etĀ al. (2023a) and Harikane etĀ al. (2023), by selecting broad-line AGN (BLAGN, or TypeI) among NIRSpec spectra (Ferruit etĀ al., 2022; Jakobsen etĀ al., 2022), have found significant AGN excess at with respect to the AGN luminosity functions derived using X-ray data (Giallongo etĀ al., 2019).
@@
The X-ray emission, produced in AGN by inverse Compton scattering of the UV photons coming from the SMBH accretion disc, has been traditionally used to select AGN. Most of the high-z AGN selected using JWST spectroscopy lie in fields covered by some of the deepest X-ray observations ever performed, like the 7Ms Chandra Deep Field South (CDFS) (Luo etĀ al., 2017), Chandra Deep Field North (CDFN) (Xue etĀ al., 2016) and AEGIS-XD (Nandra etĀ al., 2015). However, most of these newly discovered high-z AGN lack any X-ray emission (e.g. Mazzolari et al., in prep., Scholtz etĀ al., 2023a; Ćbler etĀ al., 2023; Maiolino etĀ al., 2024a; Kocevski etĀ al., 2023; Wang etĀ al., 2024; Juodžbalis etĀ al., 2024a), with only very few exceptions (Goulding etĀ al., 2023; Kovacs etĀ al., 2024; Kocevski etĀ al., 2023). This makes the picture even more difficult to interpret and suggests different scenarios such as Compton thick obscuration (Gilli etĀ al., 2022) or intrinsic X-ray weakness (Simmonds etĀ al., 2016).
@@
The difficulties in selecting high-z AGN are even more important for TypeII or narrow-line AGN (NLAGN). Different observational works have already demonstrated that some of the traditional well-known AGN emission-line diagnostic diagrams, like the BPT (Baldwin etĀ al., 1981) diagram, are no longer effective in the high-z Universe (Maiolino etĀ al., 2023a; Scholtz etĀ al., 2023a; Harikane etĀ al., 2023; Kocevski etĀ al., 2023; Ćbler etĀ al., 2023; Groves etĀ al., 2006). Indeed, at high-z, the environments are systematically more metal poor and the stellar populations are generally younger, increasing the ionization parameter () of star-forming galaxies (SFG). The latter effect, on the one hand, makes SFGs move towards the AGN locus on the BPT; on the other hand the lower metallicity of the Narrow Line Region (NLR) of AGN makes its line ratios move towards and overlap with the star formation locus on these traditional diagnostic diagrams. The detection of high-ionization emission lines (like NāVā, [NeāIV], [NeāV]) can still be considered as a safe tracer for the presence of an AGN (Calabro etĀ al., 2024; Cleri etĀ al., 2023; Brinchmann, 2023; Chisholm etĀ al., 2024), but their detection is still difficult even in the deepest JWST spectroscopic surveys like the JWST Advanced Deep Extragalactic Survey (JADES, Bunker etĀ al., 2023; Scholtz etĀ al., 2023a).
@@
However, Brinchmann (2023) first noticed that the anomalously high [OIII]emission in some of the early NIRSpec spectra at z8 could be indicating the presence of an AGN. More recently, Ćbler etĀ al. (2024) have suggested that a strong ratio between this line and the Hline could be a possible indicator for the presence of an AGN. In their work, by using JWST/NIRSpec Integral Field Spectroscopy (IFS) observations, they noticed a spatial correspondence between the AGN broad Hemission and the [OIII]line emission. They also mention the possibility that the enhanced [OIII]emission in correspondence of the AGN could be related to higher ISM temperatures driven by the AGN activity.
@@ 2.1 Low redshift samples
The sample of low-z SFGs is constituted by both normal SFGs and also by the so-called local analogues of high-z galaxies, i.e. local galaxies that show intrinsic properties similar to those observed in high-z galaxies, in terms of metallicities and emission lines ratios (e.g. Izotov etĀ al., 2018, 2019, 2021). The Sloan Digital Sky Survey (SDSS) DR7 (Abazajian etĀ al., 2009) gives the most numerous sample, mostly made of galaxies. Starting from the whole catalog provided by MPA/JHU, we considered only sources with a in all the lines involved in the diagnostic diagrams we are going to present: [OIII], [OIII], [NeāIII], [OII], and H. Then we distinguished between AGN and SFGs using the BPT diagram (Baldwin etĀ al., 1981), taking the AGN demarcation line provided by Kewley etĀ al. (2001) and after considering a further cut in S/N also in the H, Hand [NāII]Ā lines. Furthermore, we exclude from the selection AGN classified in the SDSS catalog as BLAGN, since the catalog does not provide separately the NL and BL components of the Balmer lines and our diagnostics rely on the NL emission only. We also correct all the lines for dust attenuation using the Calzetti etĀ al. (2000) attenuation law. The final SDSS SFG (AGN) sample considered in this work contains () sources.
@@
The compilation of TypeII AGN includes three sources presented in Seyfert (1943), i.e. NGC1068, NGC1275, and the core of NGC4151. These spectra were obtained at the Mount Wilson Observatory and are among the first galaxies hosting active SMBH ever observed. These three sources clearly show the [OIII]emission line in their spectra.
@@
The last sample of local AGN is taken from the MOSFIRE Deep Evolution Field (MOSDEF) survey (Kriek etĀ al., 2015), which targeted the rest-frame optical spectra of H-band-selected Galaxies at . We took the full MOSDEF catalogue and selected only those sources with a [OIII]line with a , and then we distinguished between AGN and SFG using the BPT diagram. With these S/N cuts, we could identify only 3 sources, all classified as AGN. This indicates the difficulty in detecting such a faint line further than in the local Universe in the pre-JWST era.
@@ 2.2 High redshift samples
The High-z samples of AGN and SFGs only comprise sources observed spectroscopically using JWST, the only instrument with enough resolution, sensitivity and IR-coverage to detect and disentangle the [OIII]line at . We considered the full publicly released sample of sources with a medium resolution (MR, ) spectrum in the CEERS (PID: 1345; Finkelstein etĀ al., 2022) and JADES (PID: 1210; Bunker etĀ al., 2023) surveys. For JADES, we considered the emission line fluxes and the AGN selection described in Scholtz etĀ al. (2023a). Starting from a sample of 110 sources with medium resolution (MR) spectroscopy, a reliable redshift and sufficient wavelength coverage, the authors performed a NLAGN selection using multiple AGN rest-frame UV and optical diagnostic diagrams, finally selecting 28 reliable AGN candidates in the redshift range . The diagnostic diagrams involved in the selection include the BPT (Baldwin etĀ al., 1981), the VO87 (Veilleux & Osterbrock, 1987), the CIV/CIII] vs CIII] / HeII and the detection of high-ionization emission lines such as [NeV] [NeIV], NV. To select NLAGN the authors consider in these diagnostics both demarcation lines already defined in the literature, while in other cases (such as the BPT and VOT) they defined new demarcation lines derived considering the distribution of photoionization models (Gutkin etĀ al., 2016; Feltre etĀ al., 2016; Nakajima & Maiolino, 2022). Of all the initial JADES sources, only 19 show a [OIII]detection, 7 of which AGN. We followed the same approach for the CEERS program: starting from 313 MR spectra published by the CEERS collaboration, we selected 217 sources with a reliable spectroscopic redshift and we finally selected 45 AGN candidates (Mazzolari et al., in prep). For this work, we considered 32 sources with the [OIII]line detected among the CEERS sample, 4 of which already classified as AGN. From the same programs, we also collect the BLAGN identified by Maiolino etĀ al. (2023a) and Harikane etĀ al. (2023) at , carefully checking, and in case subtracting, the broad component of the Hline.
@@
The sample of high-z sources presented in Nakajima etĀ al. (2023) includes galaxies from three different early JWST programs (CEERS, ERO, GLASS Treu etĀ al., 2022) at . We selected only the 10 sources with the [OIII]line detected, and we marked as BLAGN those that were selected based on their broad Hline in Harikane etĀ al. (2023). For the latter, we subtract the broad Hcomponent, if present.
We also include in the high-z AGN sample the BLAGN reported in Ćbler etĀ al. (2023) at , whose JWST/NIRSpec Integral Field Spectrograph (IFS) observation shows very high ionization lines and low metallicity (1/4 solar). We also include both the components of the dual AGN candidate at reported in Ćbler etĀ al. (2024). For this source, thanks to IFS spectroscopy, the authors were able to study the displacement between the position of the Hbroad line region (BLR) and the strong [OIII]emission line centroid, which were interpreted as the emissions coming from two distinct sources 620pc apart from each other. Interestingly, the authors also found an almost perfect alignment between the [OIII]emission line peak and the peak of the HBLR, suggesting a correlation between the intensity of the [OIII]line and the presence of an AGN.
We also consider the BLAGN reported in Kokorev et al. (2023) at from JWST UNCOVER Treasury survey (Labbe et al., 2021), showing a robust Hbroad component and an unprecedented black hole to host galaxy mass of at least . An even larger black hole to host galaxy mass ratio was more recently found in the BLAGN analyzed in Juodžbalis et al. (2024b), representing an extreme example of dormant SMBH at high-z. We also considered the BLAGN, triply imaged and lensed by the galaxy cluster Abell2744-QSO1, reported by Furtak et al. (2023). This source is probably undergoing a phase of rapid SMBH growth and is also heavily obscured, since the authors measured an from the Balmer decrement.
@@
We finally included in the sample also GN-z11, an exceptionally luminous galaxy at z = 10.6. Its JWST spectrum reveals the presence of an AGN through the detection of semi-forbidden lines tracing very high densities (inconsistent with the ISM but typical of the Broad Line Region), other transitions typical of AGN, fast outflows, ionization cones, and a larger Ly halo, consistent with those seen in lower redshift quasars (Maiolino etĀ al., 2024b, 2023b; Scholtz etĀ al., 2023b).
@@
| Ćbler etĀ al. (2023) | AGN | 1 | ||
| Ćbler etĀ al. (2024) | AGN | 2 | ||
| Juodžbalis et al. (2024b) | AGN | 1 | ||
@@ 3 Results
On the right panel, we show the distribution of the photoionization models presented in Sec. 2.3 according to the same line ratios. We see here that SFG models are predicted to occupy a limited region of this diagram, whose upper boundary almost perfectly corresponds with the distribution of SDSS SFG and local analogs, while AGN are expected to cover both the region occupied by SFGs and also the part of the diagnostic covered by the local AGN samples. Therefore, SFG models are not able to reach the part of the diagram occupied by local AGN, as expected also from the distribution of the observational samples. The opposite is not true, since AGN can occupy the region populated by SFGs, as observed for some local and high-z AGN. These considerations remain true even considering the full grid of SFG models computed in Feltre etĀ al. (2016), as shown in Fig. Ā 7.
@@
It is worth noting the position of some high-z sources in this diagnostic. The AGN reported in Kokorev etĀ al. (2023) shows an extremely strong [OIII]line and it is placed in a region that is not covered by any model, which probably suggests a very high electron temperature and , as also pointed out in Kokorev etĀ al. (2023). Another significant high-z AGN falling in the AGN-dominated region of this diagnostic is the source corresponding to the BLR centroid in the dual source described in Ćbler etĀ al. (2024), where, in the IFS map, the peak of the broad-line emission spatially corresponds to the peak of the [OIII]line emission. The other component, the [OIII]centroid, showing a fainter [OIII]line emission, is close to the border between the local analogues and the local AGN distribution, together with most of the high-z BLAGN. The AGN region is also populated by different AGN selected among the CEERS and JADES MR NIRSPEC spectra, but we also have different sources not selected as AGN but falling in the AGN-dominated region, as pointed out above. We demonstrate in Sec. 4 that, from the stack of their spectra, we have an indication that some of these sources may host a BLAGN too.
@@
It is worth noting that the effect of dust reddening on the line ratio [OIII]/[OII], whose components can suffer a different amount of attenuation due to their distance in wavelengths, would move data points towards the right. This can potentially shift AGNs into the SFG+AGN region under significant reddening, but not the opposite, i.e. galaxies dominated by star formation would not contaminate the AGN-only region.
@@
The diagnostic diagram presented in this section has some similarity with the so-called āOHNOā diagram that involves the [OIII]/Hvs [NeāIII]/[OII]line ratios (Backhaus etĀ al., 2022, 2023; Zeimann etĀ al., 2015; Cleri etĀ al., 2023; Trump etĀ al., 2023; Feuillet etĀ al., 2024; Killi etĀ al., 2023). The latter is mainly an ionization-sensitive diagram (ionization parameter, shape of the ionizing continuum), but it is also sensitive to metallicity and indeed at high redshift can suffer from low-metallicity galaxy contamination (Scholtz etĀ al., 2023a). The O3Hg vs Ne3O2 diagnostic is instead more stable, also at high redshift, as the separation between the AGN and SFG populations is principally based on a different gas temperature, as we will discuss further in Sec.Ā 4.
@@
The third diagnostic is presented in Fig.Ā 3 and is based on the ratio between the [OIII]/Hcompared to [OIII]/[OIII]. This diagnostic was reported also in Ćbler etĀ al. (2024). Looking at the distribution of SDSS SFGs and AGN in the left panel, we note that there is more overlap between the two populations with respect to the previous diagnostics, but there is still a region, characterized by high values of [OIII]/Hand high [OIII]/[OIII], that is populated only by local AGN. The local analogs distribution seems again to be an extrapolation of the trend of the SDSS SFG distribution. In this diagram, there are very few high-z galaxies that are in the AGN-only region of the diagnostic, while most of them are shifted towards lower values of [OIII]/[OIII], even in a region that is not covered by the current photoionization models, as it is evident from the right panel. This trend is emblematically represented by the two extreme AGN reported in Kokorev etĀ al. (2023) and in Furtak etĀ al. (2023). Given the strength of their [OIII]line emission, they are placed in the upper-left part of the plot in a region that cannot be reproduced by any model. The same trend is shown by most of the galaxies that were not selected as AGN, but that lie in the AGN region of the two diagnostics presented above. The distribution of the SFG models matches the distribution of the local galaxies, in particular in its boundary with the AGN-only region, while AGN models cover both the region occupied by SFGs and the one populated only by AGN samples.